# Dissipative solutions to Stochastic 3D Euler equations

### v2

@@ 2.4. Induction scheme
Step 2: Construction of $(u_{t\wedge\mathfrak{t}},p_{t\wedge\mathfrak{t}})$. We iteratively apply Lemma˜2.2, starting from the initial tuple $(z_{0},v_{0},p_{0},R_{0},\phi_{0})$ defined in Step 1. We obtain a sequence of dissipative Euler-Reynolds flows, $(v_{q},p_{q},R_{q},\phi_{q},z_{q})$, and by (2.18) we have that $v_{q}$ converges $\mathbb{P}$-almost surely to some $v$ in $C_{t}^{0}C^{\beta}_{x}$ for any $\beta<\alpha$. Similarly, by (2.19), $p_{q}$ converges $\mathbb{P}$-almost surely to some $p$ in $C_{t}^{0}C^{2\beta}_{x}$. Moreover, by construction $z_{q}$ converges $\mathbb{P}$-almost surely to $z$ uniformly on the time interval $0,\mathfrak{t}]$, while $\|R_{q}\|_{0}$ and $\|\phi_{q}\|_{0}$ vanish by (2.16) and (2.17). Notice that the tuple $(v,p,z)$ is $\{\mathbb{F}_{t}\}$-progressively measurable since each $(v_{q},p_{q},z_{q})$ is so, by Lemma˜2.2. Thus, we have built a solution of the system (2.5) on the time interval $0,\mathfrak{t}]$. By the discussion in Section˜2.1, we have that the process $(u,p)$ with $u:=v+z$ solves (SE) and (LEE) on the same time interval, and has the regularity stated in ˜1.1.
Step 2: Continuation of $(u,p)$. The proof is similar to [LuLuZh25, Corollary 1.8]. Let $\tilde{W}_{t}:=W_{t+\mathfrak{t}}-W_{\mathfrak{t}}$ be a cylindrical Wiener process on the filtration $\tilde{\mathbb{F}}_{t}:=\sigma(\{\tilde{W}_{s}\}_{s\leq t},u_{\mathfrak{t}},p_{\mathfrak{t}})$. We intend to apply [LuLuZh25, Theorem 1.7] with initial condtion $u^{in}:=u_{\mathfrak{t}}\in C^{\beta}_{x}$ $\mathbb{P}$-almost surely. This produces at least two distinct solutions $(\tilde{u}^{1},\tilde{p}^{1})$ and $(\tilde{u}^{2},\tilde{p}^{2})$ to (SE) with noise $Q^{1/2}d\tilde{W}$, $\mathbb{P}$-almost surely of class $\tilde{u}^{i}\in C^{0}_{t}C^{\vartheta}_{x}$ with arbitrary $\vartheta<\beta$. The pressure fields $\tilde{p}^{i}$ can be recovered from $\tilde{u}^{i}$ given by the aforementioned theorem by solving the equations $\Delta\tilde{p}^{i}=\mathord{{\rm div}}\mathord{{\rm div}}(\tilde{u}^{i}\otimes\tilde{u}^{i})$ subject to the boundary conditions $\int_{\mathbb{T}^{3}}\tilde{p}^{i}(t,x)dx=\int_{\mathbb{T}^{3}}p_{\mathfrak{t}}(x)dx$ for every $t\geq 0$, and are of class $\tilde{p}^{i}\in C_{t}C^{2\vartheta}_{x}$ $\mathbb{P}$-almost surely by [CoDR20, Corollary 3.3].
Then the glued processes $(u^{i}_{t},p^{i}_{t}):=(u_{t},p_{t})\mathbf{1}_{\{t\leq\mathfrak{t}\}}+(\tilde{u}^{i}_{t-\mathfrak{t}},\tilde{p}^{i}_{t-\mathfrak{t}})\mathbf{1}_{\{t>\mathfrak{t}\}}$, $i=1,2$, are solutions to (SE) that satisfy (LEE) by construction on the time interval $t\in[0,\mathfrak{t}]$. Since $\vartheta<\beta<\alpha$ is arbitrary, each $(u^{i},p^{i})$ has the space regularity stated in ˜1.1, concluding the proof.
@@
(2.25) $\|u_{q}(t)\|_{N}+\|v_{q}(t)\|_{N}\lesssim\begin{cases}\lambda_{q}^{N}\delta_{q}^{1/2},\quad&N>0\\ 1\,,\quad&N=0\end{cases},$
where the implicit constant does not depend on the parameter $a$. This independence of $a$ will be crucial in the following sections. Indeed, we will be able to absorb various implicit constants emerging from the convex integration scheme by picking $a$ large enough. Since this remark is not trivial just for $N=0$, let us focus on that case. Indeed, since $a>2$ it holds for every $q\in\mathbb{N}$

### v3

@@ 2.3. Stochastic Euler Equations: Induction scheme
Step 2: Construction of $(u_{t\wedge\mathfrak{t}},p_{t\wedge\mathfrak{t}})$. We iteratively apply Lemma 2.1, starting from the initial tuple $(z_{0},v_{0},p_{0},R_{0},\phi_{0})$ defined in Step 1. We obtain a sequence of dissipative Euler-Reynolds flows, $(v_{q},p_{q},R_{q},\phi_{q},z_{q})$, and by (2.18) we have that $v_{q}$ converges $\mathbb{P}$-almost surely to some $v$ in $C_{t}^{0}C^{\beta}_{x}$ for any $\beta<\alpha$. Similarly, by (2.19), $p_{q}$ converges $\mathbb{P}$-almost surely to some $p$ in $C_{t}^{0}C^{2\beta}_{x}$. Moreover, by construction $z_{q}$ converges $\mathbb{P}$-almost surely to $z$ uniformly on the time interval $0,\mathfrak{t}]$, while $\|R_{q}\|_{0}$ and $\|\phi_{q}\|_{0}$ vanish by (2.16) and (2.17). Notice that the tuple $(v,p,z)$ is $\{\mathbb{F}_{t}\}$-progressively measurable since each $(v_{q},p_{q},z_{q})$ is so, by Lemma 2.1. Thus, we have built a solution of the system (2.8) on the time interval $0,\mathfrak{t}]$. By the discussion in subsection 2.3, we have that the process $(u,p)$ with $u:=v+z$ solves (SE) and (LEE) on the same time interval, and has the regularity stated in Theorem 1.1.
Step 3: Continuation of $(u,p)$. The proof is similar to [LuLuZh25, Corollary 1.8]. Let $\tilde{W}_{t}:=W_{t+\mathfrak{t}}-W_{\mathfrak{t}}$ be a cylindrical Wiener process on the filtration $\tilde{\mathbb{F}}_{t}:=\sigma(\{\tilde{W}_{s}\}_{s\leq t},u_{\mathfrak{t}},p_{\mathfrak{t}})$. We intend to apply [LuLuZh25, Theorem 1.7] with initial condtion $u^{in}:=u_{\mathfrak{t}}\in C^{\beta}_{x}$ $\mathbb{P}$-almost surely. This produces at least two distinct solutions $(\tilde{u}^{1},\tilde{p}^{1})$ and $(\tilde{u}^{2},\tilde{p}^{2})$ to (SE) with noise $Q^{1/2}d\tilde{W}$, $\mathbb{P}$-almost surely of class $\tilde{u}^{i}\in C^{0}_{t}C^{\vartheta}_{x}$ with arbitrary $\vartheta<\beta$. The pressure fields $\tilde{p}^{i}$ can be recovered from $\tilde{u}^{i}$ given by the aforementioned theorem by solving the equations $\Delta\tilde{p}^{i}=\mathord{{\rm div}}\mathord{{\rm div}}(\tilde{u}^{i}\otimes\tilde{u}^{i})$ subject to the boundary conditions $\int_{\mathbb{T}^{3}}\tilde{p}^{i}(t,x)dx=\int_{\mathbb{T}^{3}}p_{\mathfrak{t}}(x)dx$ for every $t\geq 0$, and are of class $\tilde{p}^{i}\in C_{t}C^{2\vartheta}_{x}$ $\mathbb{P}$-almost surely by [CoDR20, Corollary 3.3].
Then the glued processes $(u^{i}_{t},p^{i}_{t}):=(u_{t},p_{t})\mathbf{1}_{\{t\leq\mathfrak{t}\}}+(\tilde{u}^{i}_{t-\mathfrak{t}},\tilde{p}^{i}_{t-\mathfrak{t}})\mathbf{1}_{\{t>\mathfrak{t}\}}$, $i=1,2$, are solutions to (SE) that satisfy (LEE) by construction on the time interval $t\in[0,\mathfrak{t}]$. Since $\vartheta<\beta<\alpha$ is arbitrary, each $(u^{i},p^{i})$ has the space regularity stated in Theorem 1.1, concluding the proof.
@@
(2.29) $\|u_{q}(t)\|_{N}+\|v_{q}(t)\|_{N}\lesssim\begin{cases}1\,,\quad&N=0,\\ \lambda_{q}^{N}\delta_{q}^{1/2},\quad&N\geq 1,\end{cases}$
where the prefactor does not depend on the parameter $a$. This independence of $a$ will be crucial in the following sections. Indeed, we will be able to absorb various implicit constants emerging from the convex integration scheme by picking $a$ large enough. Since this remark is not trivial just for $N=0$, let us focus on that case. Indeed, since $a>2$ it holds for every $q\in\mathbb{N}$